Please give me the answer to: 5% interest on $6,000.00 for 9 years...................Thank you
Answer by rfer(10432)

FV=P(1+r)^t
FV=6000(1.05)^9
FV=6000(1.551328216)
FV=$9307.97
interest earned=$3307.97